OpenStudy (anonymous):

Find the diameter, to the nearest mm, of a sphere with surface area 215.8cm squared.

OpenStudy (anonymous):

\[surface\:\:\:\:area =4* \pi*r^2\] \[ 215.8 = 4*\pi*r^2\] \[2r =d\] r is radius and d is diameter

OpenStudy (anonymous):

i feel it is 4.143176208 cm

OpenStudy (anonymous):

I have the answer with me but it is not the one above. I am confused on how to solve this problem.

OpenStudy (anonymous):

what is the formula of sphere surface area?

OpenStudy (anonymous):

SA=4piersquared

OpenStudy (anonymous):

thus, 215.8 =4*(22/7)*r^2 r^2 = 215.8/(4*3.14) r = (17.16590909)^(1/2) r =4.143176208 cm this is ur answer now u can cross check it!!!!!!!!!!!

OpenStudy (anonymous):

The answer in the back of book says the diameter is 83mm

OpenStudy (anonymous):

ok buddies now feeling lil bit tired now goin to have a cup tea.................

OpenStudy (anonymous):

will be back soon............

OpenStudy (anonymous):

enjoy your tea!

OpenStudy (anonymous):

I did it again it's alright maybe some misprint in your book

OpenStudy (anonymous):

Okay, I guess so. Thanks for the help though.

OpenStudy (anonymous):

No Problem

OpenStudy (anonymous):

hey ur answer is correct diameter = 2 * radius Diameter = 2* 4.143176208 =8.286352416cm aprox. 83 mm

OpenStudy (anonymous):

Thanks Sheg!

OpenStudy (anonymous):

This posting corrects an earlier posting which I deleted. To convert cm^2 to mm^2 the decimal point has to be shifted two places to the right (multiplyied by 100). Solve the following for d:\[21580=4 \pi \left(\frac{d}{2}\right)^2\]d = 82.8802 or 83 mm^ rounded to the nearest mm.